OpenStudy (anonymous):

Need help, -22fg -4f -9fg -5f = 31fg - (-1f) Is that correct?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (danjs):

you mean, is the equality true as it is like that?

OpenStudy (danjs):

for a bunch of terms added or subtracted like that, you can combine together the ones with the same variables, looks like you got that , the number in front totals , -22fg -4f -9fg -5f = (-22 - 9)*fg + (-4 -5)*f or -31fg - 9f

OpenStudy (anonymous):

@DanJS -22fg - 4f - 9f - 5f= -22fg + -9fg - 4f - 5f= 31fg - (4f - 5f)= 31fg - (-1f) That's how I got it

OpenStudy (anonymous):

9fg*

OpenStudy (anonymous):

I did the subtraction rule for the second part and put the like terms together

OpenStudy (danjs):

i see, -22 + (-9) = -22 - 9 = -31 for the fg term not sure if you may just forgot the - sign

OpenStudy (danjs):

-4 - 5, you cant take the negative from just one of them and leave the other the same, if you want to pull out a -1 from both first you can -4 - 5 = (-1)*( 4 + 5 ) = -1 * 9 = -9

OpenStudy (anonymous):

so am I right?

OpenStudy (danjs):

you had a -1*(4f - 5f), you changed the sign on the 5 to a plus term, = -4f + 5f in that case

OpenStudy (anonymous):

it isn't -4f, that's a subtraction sign

OpenStudy (danjs):

to make it simpler, just always add things, and if the number is negative use a negative , like ... +( -4f - 5f) = ...+ (-9f)

OpenStudy (anonymous):

so when there's a problem like that, that has subtraction signs, I always add?

OpenStudy (danjs):

subtraction is the same as addition a - b = a + (-b) addition of negative numbers

OpenStudy (danjs):

it is up to you, just total the numbers you need, if the term is written as subtraction, that is a negative number on that term

OpenStudy (danjs):

think of the number line, if you have a -22 and a -9 you move left -22 then left again -9, ending at -31 you added -9 to -22

OpenStudy (danjs):

same for -4 and -5 left 4, then left 5, ends at -9

OpenStudy (anonymous):

@DanJS But since when was 4 and 5 negatives?

OpenStudy (danjs):

in the problem Need help, -22fg -4f -9fg -5f all 4 terms are negative

OpenStudy (anonymous):

no, only -22fg is a negative, the rest are positives @DanJS

OpenStudy (anonymous):

@Fuzzbro

OpenStudy (anonymous):

alright lemme catch up.

OpenStudy (anonymous):

Are you trying to simplify the equation?

OpenStudy (anonymous):

Well I want to make sure my answer is correct

OpenStudy (anonymous):

I believe it ends up being -31fg -(-9f)

OpenStudy (anonymous):

how did you get 9?

OpenStudy (anonymous):

Because 4+5 equals 9. So -4 plus -5 equals -9

OpenStudy (anonymous):

But 4f-5f= -1 , right?

OpenStudy (anonymous):

it's the 4 isn't positive, it's negative

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

-4f -9fg -5f

OpenStudy (anonymous):

that's how the question was typed :P

OpenStudy (anonymous):

the 4 and 5 are negative

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

there

OpenStudy (anonymous):

yup the 4 and 5 are both negative

OpenStudy (anonymous):

But in k12 they would've put parenthesis right??

OpenStudy (anonymous):

they'd make it clear i'm pretty sure

OpenStudy (anonymous):

separate the fs and fgs

OpenStudy (anonymous):

-4f -5f -22fg -9fg

OpenStudy (anonymous):

it's the same deal as with the fgs

OpenStudy (anonymous):

so they're all negatives?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

well then how do I solve it?? I'm so confused

OpenStudy (anonymous):

I'm pretty sure they're not all negatives, they'ed put commas in-between the numbers

OpenStudy (anonymous):

whats -4 -5

OpenStudy (anonymous):

if that were the case

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

there is -9f now what's -22 -9

OpenStudy (anonymous):

-22 -9?

OpenStudy (anonymous):

subtracting sign?

OpenStudy (anonymous):

yeah negative twenty two, minus 9

OpenStudy (anonymous):

-31?

OpenStudy (anonymous):

Yup! so now you have -31fg -9f

OpenStudy (anonymous):

just so you know a negative sign is just the same as a minus sign

OpenStudy (anonymous):

subtracting any number from another number is like adding negatives to it

OpenStudy (anonymous):

but double negatives, or subtracting with a negative sign too is like taking away negatives

OpenStudy (anonymous):

thinking about the equation as a number line can help too

OpenStudy (anonymous):

can u give me a example of when negatives are like subtraction signs?

OpenStudy (anonymous):

lets say you have -3 and 7

OpenStudy (anonymous):

the three is negative, if you added -3 to 7 what would it be

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

exactly, subtracting is like adding a negative number to the number

OpenStudy (anonymous):

oh, I see. So are you positive it's -31fg-9f?

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

ok I hope I get a good score, brb

OpenStudy (anonymous):

Combine like terms. 8 + 16r – 9r A. 8 + 7r B. 15r C. 15 D. 8 – 7r Wait, is this A?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

basically that question is 0.825 -0.5

OpenStudy (anonymous):

so it would be 0.325x

OpenStudy (anonymous):

you're correct

OpenStudy (anonymous):

alright!

OpenStudy (anonymous):

is that the end of the test?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Got 100%!

OpenStudy (anonymous):

Thank you so muchhh!

OpenStudy (anonymous):

Yay :D

OpenStudy (anonymous):

no problem at all :D

OpenStudy (anonymous):

Ok i'm going to close the question

OpenStudy (anonymous):

alright message me if you need anything

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i gtg for now though

OpenStudy (anonymous):

oh bye~!